# GeoLLaVA-8K: Scaling Remote-Sensing Multimodal Large Language Models to 8K Resolution

### v1

@@ 6 Conclusion
In this paper, we addressed two fundamental challenges in scaling vision-language models to UHR RS imagery: the lack of suitable training data and the computational burden of token explosion. To tackle these issues, we introduced two new UHR image-text datasets, SuperRS-VQA and HighRS-VQA, which greatly expand the data available for vision-language tasks on UHR RS images. We also proposed two novel token-efficient strategies: Background Token Pruning and Anchored Token Selection, which effectively reduce the number of visual tokens processed from 8K-resolution images while preserving essential information. Building on these contributions, we developed GeoLLaVA-8K, the first RS-specific MLLM that can directly handle inputs up to 8K×8K resolution. GeoLLaVA-8K achieved state-of-the-art performance on the XLRS-Bench benchmark, outperforming both open- and closed-source MLLMs and demonstrating the effectiveness of our token-efficient approach. These results underscore the value of domain-adapted, token-efficient modeling and provide a new foundation for high-fidelity image understanding in RS. Limitation. We have primarily focused on optical satellite imagery; evaluating GeoLLaVA-8K on other sensor modalities (e.g., synthetic aperture radar or multispectral images) will be important to ensure broader applicability.
@@ A.2 More Details of GeoLLaVA-8K
Tab. 5 presents the key training configuration of our GeoLLaVA-8K model.
The model is trained on 81K UHR geospatial image-text pairs. We use different learning rates for the visual components (1e-6) and the projection layers interacting with the LLM (5e-6), and optimize the model using ZeRO-2 parallelism with a batch size of 32 for one training epoch.
@@ A.3 Sub-tasks (L-2 capability) Results on GeoLLaVA-8K
Sub-tasks (L-2 Capability) Counting Scene Classification Object Spatial Relationship Object Properties Planning Anomaly Reasoning Complex Reasoning Spatiotemporal Reasoning
@@ 16   OOM
As shown in Tab. 9, although VRSBench contains approximately ten times more samples than SuperRS-VQA, training solely on SuperRS-VQA yields significantly better performance (49.4 vs. 34.8). This is likely attributed to the substantial resolution gap: VRSBench averages 512$\times$512 pixels, while SuperRS-VQA averages 8,376$\times$8,376, and also reflects the higher annotation quality of SuperRS-VQA. A further accuracy boost to 51.4 was observed when incorporating the HighRS-VQA dataset, validating its complementary value and effectiveness.
@@
Figure 9: Visual tokens and their semantics detected by the $logit$ $lens$ technique.
@@ A.7.1 Motivation
    A: Ultra-high-resolution (UHR) remote sensing (RS) imagery offers valuable data for Earth observation but pose challenges for existing multimodal foundation models due to the key bottlenecks: limited availability of UHR training data. To address data scarcity, we introduce SuperRS-VQA (avg. 8,376$\times$8,376) and HighRS-VQA (avg. 2,000$\times$1,912), the highest-resolution vision-language datasets in RS to date, covering 22 real-world dialogue tasks.

### v2

@@ 6 Conclusion
In this paper, we addressed two fundamental challenges in scaling vision-language models to UHR RS imagery: the lack of suitable training data and the computational burden of token explosion. To tackle these issues, we introduced two new UHR image-text datasets, SuperRS-VQA and HighRS-VQA, which greatly expand the data available for vision-language tasks on UHR RS images. We also proposed two novel token-efficient strategies: Background Token Pruning and Anchored Token Selection, which effectively reduce the number of visual tokens processed from 8K-resolution images while preserving essential information. Building on these contributions, we developed GeoLLaVA-8K, the first RS-specific MLLM that can directly handle inputs up to $\times8K$ resolution. GeoLLaVA-8K achieved state-of-the-art performance on the XLRS-Bench benchmark, outperforming both open- and closed-source MLLMs and demonstrating the effectiveness of our token-efficient approach. These results underscore the value of domain-adapted, token-efficient modeling and provide a new foundation for high-fidelity image understanding in RS.
Limitation. We have primarily focused on optical satellite imagery; evaluating GeoLLaVA-8K on other sensor modalities (e.g., synthetic aperture radar or multispectral images) will be important to ensure broader applicability.
Acknowledgements: This work was partially supported by the National Natural Science Foundation of China (No. 62372459, No.62376282, No.62225113 and No. 624B2109).
@@ A.2 More Details of GeoLLaVA-8K
Tab. 9 presents the key training configuration of our GeoLLaVA-8K model. The model is trained on 81K UHR geospatial image-text pairs. We use different learning rates for the visual components (1e-6) and the projection layers interacting with the LLM (5e-6), and optimize the model using ZeRO-2 parallelism with a batch size of 32 for one training epoch.
@@ A.3 Sub-tasks (L-2 capability) Results on GeoLLaVA-8K
Sub-tasks (L-2 Capability)
Counting
Scene
Classification
Object
Spatial Relationship
Object
Properties
Planning
Anomaly
Reasoning
Complex
Reasoning
Spatiotemporal
Reasoning
@@ 16   OOM
As shown in Tab. 13, although VRSBench contains approximately ten times more samples than SuperRS-VQA, training solely on SuperRS-VQA yields significantly better performance (49.4 vs. 34.8). This is likely attributed to the substantial resolution gap: VRSBench averages $\times512$ pixels, while SuperRS-VQA averages $\times8,376$, and also reflects the higher annotation quality of SuperRS-VQA. A further accuracy boost to 51.4 was observed when incorporating the HighRS-VQA dataset, validating its complementary value and effectiveness.
@@
Figure 9: Visual tokens and their semantics detected by the logitlogit lenslens technique.
@@ A.7.1 Motivation
[…]